OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 3y = 6 x = -three halvesy + 3 3y = -2x + 6 y = -two thirdsx + 2

OpenStudy (anonymous):

y = -two thirdsx + 2 Because Slope intercept corm should look like y=mx+b

OpenStudy (31356):

2x+3y=6 because it is in slop-intercept form, which is y=mx+b! There you go.

OpenStudy (anonymous):

d

OpenStudy (31356):

I don't think it is b

OpenStudy (31356):

Actually d

OpenStudy (anonymous):

its not b

OpenStudy (anonymous):

D is right

OpenStudy (31356):

Never mind I was wrong it is D

OpenStudy (31356):

Sorry, I overlooked.

OpenStudy (anonymous):

lol